A hawk can fly 300 miles in 8 hours with the wind. Flying against the wind, the hawk covers only one-third of the distance in 7 hours. What is the rate of the wind?
Answer by mananth(16946)

speed of hawk in still air = x
speed of wind = y
distance / speed = time
300/(x+y)=8
8x+8y=300
...
against wind
100/(x-y)=7
7x-7y=100
...
multiply first equation by 7 and second equation by 8 and add them
56x+56y=2100
56x-56y=800
...
112x=2900
x= 25.89
plug the value of x in any equation you will get y = 11.6 speed of wind